# A Joint X-ray and Optical Study of the Massive Redback Pulsar J2215+5135

### v1

@@ 4.2 X-ray Models
Figure 4: The best-fit IBS X-ray light curves. One model fit fixes $i$ at the best-fit value from the optical modeling, while the other model fit leaves $i$ as a free parameter.
@@
We fit for $\beta$, $f_{v}$, and $k$. We perform fits with fixed $i=71^{\circ}$, the global minimum value from the optical fits, and with $i$ left as a free parameter. We fix $\Gamma_{nose}=1.1$, consistent with numerical models (Bogovalov et al., 2008; Dubus et al., 2015; Kandel et al., 2019). At the IBS nose, we set $B_{IBS}=3B_{LC}\left(R_{LC}/r_{0}\right)\approx 70$ G (where $B_{LC}$ is the magnetic field at the light cylinder radius $R_{LC}$; the factor of 3 is due to relativistic shock compression; Sullivan & Romani, 2023), based on the inferred orbital separation $a\approx 1.5\times 10^{11}$ cm, $P_{s}=2.61$ and $\dot{P}_{s}=2.38\times 10^{-20}$ (Smith et al., 2023). We set $p=0.76$ so that our IBS model spectrum has $\Gamma_{IBS}=0.94$ in the $0.5-10$ keV band, consistent with the actual spectral results shown in Table 1. Our fits prefer a forward phase shift $\delta\Phi\sim 0.03$ from the radio/gamma-ray determined ephemeris and from our optical phase. We leave $\delta\Phi$ as a fit parameter. We also fit the injected particle spectrum normalization $\dot{N}_{0}$ and a phase independent flux $F_{b}$ added as an offset. We add the offset to account for additional X-ray flux present in the source, possibly due to a pulsar wind nebula (PWN) or the magnetosphere; this component is also present in the spectral fits, as noted above.
@@
Our best-fit light curves are shown in Fig. 4 and the parameter estimates are shown in Table 3. Our model fits the data moderately well, with $\chi^{2}/$DoF$=1.45$ when fixing $i$ and $\chi^{2}/\text{DoF}=1.16$ when leaving $i$ free. This corresponds to an Akaike information criterion (AIC) value of 41 for the fixed $i$ model and 36 for the free $i$ model. Departures from a simple light curve in the phase range $\Phi=0.6-0.7$ contribute a large factor to the $\chi^{2}$. In examining the orbit-to-orbit variations, we find that this phase range has somewhat higher variability than the rest of the light curve, but we are unable to attribute the structure in this phase range to IBS flares (as have been seen in other spider pulsars; e.g. An et al., 2017). Thus, the flux dip at $\Phi\approx 0.6$ might represent an absorption component, or, alternatively, the flux increase at $\Phi\approx 0.65$ might represent an extra emission component, possibly structure in a thick post-shock flow not captured by our thin-shock geometry. We have left these points in the fit to be conservative; excising these points barely changes the fixed $i$ best-fit parameters. We also note the low bin at $\Phi=0.25$. While at the appropriate phase for a companion eclipse, its significance is low. Furthermore, we would not expect significant occultation of the extended IBS at our moderate $\sim 70^{\circ}$ inclination.
@@
The model fitting gives values $\beta\approx 2$ and $f_{v}\approx 3-6$, signaling a rather flat IBS and slow companion wind. If left free, $i$ decreases to $52^{+17}_{-9}$ deg to accommodate the bridge flux (including the $\Phi\approx 0.65$ component). Such low $i$ implies $M_{NS}=3.8^{+3.9}_{-1.4}$ M; the central value is unphysical, although less than $2\sigma$ different from our optical $i$ mass determination. Indeed, $i$ itself is only in $\sim 1\sigma$ disagreement with the optical value. We thus consider the fixed $i$ IBS fit more reliable.
@@ 5 Discussion and Conclusion
Our optical and X-ray model fits, while imperfect, provide useful estimates of the parameters of J2215. Having used the radial velocity $K_{c}$ from Kandel & Romani (2020), we believe that these new values should supersede the results of that paper and those in the less realistic modeling of Linares et al. (2018). Our revised value for the mass is just about $1\sigma$ less than the Kandel & Romani (2020) estimate, but, at $M_{NS}=2.15\pm 0.10$ M, remains more than $1\sigma$ larger than 2 M.
With well-detected light curve peaks, our X-ray results are much more robust than those of Romani & Sanchez (2016). Our modeling suggests $\beta\lesssim 2$ and $f_{v}<10$. From these results and $\dot{E}=5.2\times 10^{34}$ $I_{45}$ erg s-1 (Smith et al., 2023), we estimate the companion mass-loss rate
@@
The estimated value of the phase-independent flux from the light curve fits $F_{b}\approx 5\times 10^{-14}$ erg cm-2 s-1 is in good agreement with the phase-independent flux estimated from the spectral fitting $F_{0}$. These fluxes correspond to a luminosity of $7\times 10^{31}$ erg s-1 at $d=3.3$ kpc and could originate from magnetospheric emission or, more likely, from a faint, compact PWN ( e.g. Li et al., 2008). The spectral index $\Gamma_{0}\approx 1.5$ is quite consistent with a PWN source, while the luminosity of $10^{-3}{\dot{E}}$ is similar to other observed PWNe efficiencies. The slightly larger value of $F_{0}$ would include a small amount of emission from the wings of the IBS peaks in the off-phase.
It is also useful to compare our measured energy fluxes with the pulsar spin-down power budget ${\dot{E}}=5\times 10^{34}$ $I_{45}$ ${\rm erg\,s^{-1}}$. Several neutron star masses have been securely measured above $2M_{\odot}$, pointing towards a quite stiff equation of state. We can estimate $I_{45}\approx[0.8-1.2](M/M_{\odot})^{3/2}$ (Lattimer & Schutz, 2005), i.e. $2.5-3.8$ for $2.15$ $M_{\odot}$. The ${\rm sin}^{2}\theta_{\ast}$-distributed heating luminosity $L_{H}\approx 2.5\times 10^{34}$ ${\rm erg\,s^{-1}}$ is in good accord with the isotropic GeV gamma-ray luminosity measured by the Fermi LAT, $L_{\gamma}=3.3\times 10^{34}$ ${\rm erg\,s^{-1}}$, suggesting that our ‘direct heating’ is dominated by the GeV photons. This requires that the radiation be directed toward the companion at the pulsar spin equator, as expected from the spin-orbit aligned recycled pulsar outer magnetosphere/wind emission model (Philippov & Spitkovsky, 2018). This gamma-ray direct heating power is a substantial fraction of the spin-down energy budget, but can be comfortably accommodated for $I_{45}=2-3$. This also reduces our estimated evaporation timescale below 1 Gyr.
@@
The observed 0.5-10 keV X-ray spectrum comes from $10^{4}<\gamma<10^{5}$ $e^{\pm}$ particles. Our fit $\dot{N}_{0}\sim 3-4\times 10^{31}$ s-1 thus gives a minimum $\dot{E}_{pw}\gtrsim 5\times 10^{31}$  erg s-1 for $p\approx 0.8$. This is well below the available power ${\dot{E}}-L_{H}\approx 12\times 10^{34}$ ${\rm erg\,s^{-1}}$ for $I_{45}\approx 3$. To avoid saturating the residual spin-down power, there is an upper limit of $\gamma_{\rm max}\sim 5.5\times 10^{7}$. Higher energy X-ray and gamma-ray measurements can further probe the high $\gamma$ $e^{\pm}$ population (which cools rapidly). If magnetic reconnection is the primary acceleration mechanism (as suggested by the very hard observed spectra), the maximum electron energy should be $\gamma_{\rm max}\sim\sigma\gamma_{\rm min}$, where $\sigma=B^{2}/(4\pi\gamma_{\rm min}n_{0}m_{e}c^{2})$ is the pulsar wind magnetization and $n_{0}$ is the e± number density (Sironi & Spitkovsky, 2011, 2014). As expected, the J2215 pulsar wind should be very strongly magnetized with $10\lesssim\sigma\lesssim 6\times 10^{7}$. Finally, the companion hot spot indicates extra localized heating. Sanchez & Romani (2017) attribute companion hot spots to IBS-energized particles precipitating to the companion magnetic poles. In our fits, the small $\beta$ allows the relatively flat IBS shock to capture $\sim 40-50\%$ of the $\dot{E}_{pw}$ flux, while the thermal emission of the hot spot represents $1\times 10^{32}$ ${\rm erg\,s^{-1}}$. Only a small fraction $\sim 2\times 10^{-3}$ of the maximum available IBS-processed pulsar wind power needs to reach the surface to heat the companion magnetic pole. Simultaneous optical/UV observations would be useful to re-check the U band normalization and further constrain the hot spot contribution.
J2215 remains one of the most massive neutron stars known. Its high inferred mass-loss rate makes this system a plausible isolated MSP progenitor, while its unusually flat IBS may reprocess a large fraction of the spin-down power. The hard X-ray spectrum and sharp peaks provide an important probe of relativistic shock dynamics and particle acceleration. Higher energy observations (e.g. with NuSTAR) should further constrain the IBS $e^{\pm}$ population. Since the prominent IBS peaks indicate beamed emission and thus moderate bulk Lorentz factors in the IBS flow, the various orbital phases probe different portions of the shock with varying obliquity. Thus, IBS-dominated spider pulsars, like J2215, show emission from a range of relativistic shock geometries. Future phase-resolved spectrum and polarization measurements of these natural laboratories can be compared with numerical simulations to advance our understanding of oblique, strongly magnetized relativistic shocks.

### v2

@@ 4.2 X-ray Models
Figure 5: The best-fit IBS X-ray light curves. One model fit fixes $i=80^{\circ}$, compatible with the optical modeling, while the other model fit leaves $i$ as a free parameter.
@@
We fit for $\beta$, $f_{v}$, and $k$. We perform fits with fixed $i=80^{\circ}$, compatible with the best fitting $(\log Z,\alpha)=(-0.5,0)$ optical model, and with $i$ left as a free parameter. We fix $\Gamma_{nose}=1.1$, consistent with numerical models (Bogovalov et al., 2008; Dubus et al., 2015; Kandel et al., 2019). At the IBS nose, we set $B_{IBS}=3B_{LC}\left(R_{LC}/r_{0}\right)\approx 70$ G (where $B_{LC}$ is the magnetic field at the light cylinder radius $R_{LC}$; the factor of 3 is due to relativistic shock compression; Sullivan & Romani, 2023), based on the inferred orbital separation $a\approx 1.5\times 10^{11}$ cm, $P_{s}=2.61$ ms and $\dot{P}_{s}=2.38\times 10^{-20}$ (Smith et al., 2023). We set $p=0.76$ so that our IBS model spectrum has $\Gamma_{IBS}=0.94$ in the $0.5-10$ keV band, consistent with the actual spectral results shown in Table 1. Our fits prefer a forward phase shift $\delta\Phi\sim 0.03$ from the radio/gamma-ray determined ephemeris and from our optical phase. We leave $\delta\Phi$ as a fit parameter. We also fit the injected particle spectrum normalization $\dot{N}_{0}$ and a phase independent flux $F_{b}$ added as an offset. We add the offset to account for additional X-ray flux present in the source, possibly due to a pulsar wind nebula (PWN) or the magnetosphere; this component is also present in the spectral fits, as noted above.
@@
Our best-fit light curves are shown in Fig. 5 and the parameter estimates are shown in Table 3. Our model fits the data moderately well, with $\chi^{2}/$DoF$=1.35$ when fixing $i$ and $\chi^{2}/\text{DoF}=1.16$ when leaving $i$ free. This corresponds to an Akaike information criterion (AIC) value of 39 for the fixed $i$ model and 36 for the free $i$ model. Departures from a simple light curve in the phase range $\Phi=0.6-0.7$ contribute a large factor to the $\chi^{2}$. In examining the orbit-to-orbit variations, we find that this phase range has somewhat higher variability than the rest of the light curve, but we are unable to attribute the structure in this phase range to IBS flares (as have been seen in other spider pulsars; e.g. An et al., 2017). Thus, the flux dip at $\Phi\approx 0.6$ might represent an absorption component, or, alternatively, the flux increase at $\Phi\approx 0.65$ might represent an extra emission component, possibly structure in a thick post-shock flow not captured by our thin-shock geometry. We have left these points in the fit to be conservative; excising these points barely changes the fixed $i$ best-fit parameters. We also note the low bin at $\Phi=0.25$. While at the appropriate phase for a companion eclipse, its significance is low. Furthermore, we would not expect significant occultation of the extended IBS at $i\lesssim 85^{\circ}$.
@@
The model fitting gives values $\beta\approx 2$ and $f_{v}\approx 3-6$, signaling a rather flat IBS and slow companion wind. If left free, $i$ decreases to $52^{+17}_{-9}$ deg to accommodate the bridge flux (including the $\Phi\approx 0.65$ component). Such low $i$ implies $M_{NS}=3.8^{+3.9}_{-1.4}$ M; the central value is unphysical, although less than $2\sigma$ different from our optical mass determinations. Indeed, $i$ itself is only in $\sim 1.5\sigma$ disagreement with the possible optical values. We thus consider the fixed $i$ IBS fit more reliable.
@@ 5 Discussion and Conclusion
Our optical and X-ray model fits, while imperfect, provide more realistic estimates of the parameters of J2215. Having used the radial velocity $K_{c}$ from Kandel & Romani (2020), we believe that these new values should supersede the results of that paper and those in the less physical modeling of Linares et al. (2018), especially given the unusual sensitivity to the metallicity. Our best-fit $(\log Z,\alpha)=(-0.5,0)$ result $M_{NS}=1.98\pm 0.08$ M is more than $2\sigma$ less than the KR20 estimate. If we adopt $\log Z=0$, as suggested by the optical spectroscopy, the mass lowers by an additional $\sim 1\sigma$ to $M_{NS}=1.91\pm 0.06$ M. Overall, including the metallicity-induced systematic uncertainty, masses from $\approx 1.85-2.3$  M remain acceptable. Improved absolute photometry, especially in UV bands could help break the metallicity degeneracy. Our present result does not strongly constrain the dense matter equation of state given these uncertainties, but can still be valuable for high mass neutron star population studies (e.g. Brandes et al., 2023; Komoltsev et al., 2024).
With well-detected light curve peaks, our X-ray results are much more robust than those of Romani & Sanchez (2016). Our modeling suggests $\beta\lesssim 2.5$ and $f_{v}<10$. From these results and $\dot{E}=5.2\times 10^{34}$ $I_{45}$ erg s-1 (Smith et al., 2023), we estimate the companion mass-loss rate
@@
The estimated value of the phase-independent flux from the light curve fits $F_{b}\approx 5\times 10^{-14}$ erg cm-2 s-1 is in good agreement with the phase-independent flux estimated from the spectral fitting $F_{0}$. These fluxes correspond to a luminosity of $7\times 10^{31}$ erg s-1 at $d=3.3$ kpc and could originate from magnetospheric emission or, more likely, from a faint, compact PWN (e.g. Li et al., 2008). The spectral index $\Gamma_{0}\approx 1.5$ is quite consistent with a PWN source, while the efficiency $L\approx 10^{-3}{\dot{E}}$ is similar to that of other observed PWNe. The slightly larger value of $F_{0}$ would include a small amount of emission from the wings of the IBS peaks in the off-phase.
It is also useful to compare our measured energy fluxes with the pulsar spin-down power budget ${\dot{E}}=5\times 10^{34}$ $I_{45}$ ${\rm erg\,s^{-1}}$. Several neutron star masses have been securely measured above $2$ M, pointing towards a quite stiff equation of state. We can estimate $I_{45}\approx[0.8-1.2](M/M_{\odot})^{3/2}$ (Lattimer & Schutz, 2005), i.e. $2.2-3.4$ for $2$ $M_{\odot}$. The ${\rm sin}^{2}\theta_{\ast}$-distributed heating luminosity $L_{H}\approx 2.5\times 10^{34}$ ${\rm erg\,s^{-1}}$ is in good accord with the isotropic GeV gamma-ray luminosity measured by the Fermi LAT, $L_{\gamma}=3.3\times 10^{34}$ ${\rm erg\,s^{-1}}$, suggesting that our ‘direct heating’ is dominated by the GeV photons. This requires that the radiation be directed toward the companion at the pulsar spin equator, as expected from outer magnetosphere/wind emission models (e.g. Philippov & Spitkovsky, 2018) for spin-orbit aligned recycled pulsars. This gamma-ray direct heating power is a substantial fraction of the spin-down energy budget, but can be comfortably accommodated for $I_{45}=2-3$. This also reduces our estimated evaporation timescale below 1 Gyr.
@@
The observed 0.5-10 keV X-ray spectrum comes from $10^{4}<\gamma<10^{5}$ $e^{\pm}$ particles. Our fit $\dot{N}_{0}\sim 3-4\times 10^{31}$ s-1 thus gives a minimum $\dot{E}_{pw}\gtrsim 5\times 10^{31}$  erg s-1 for $p\approx 0.8$. This is well below the available power ${\dot{E}}-L_{H}\approx 12\times 10^{34}$ ${\rm erg\,s^{-1}}$ for $I_{45}\approx 3$. To avoid saturating the residual spin-down power, there is an upper limit of $\gamma_{\rm max}\sim 5.5\times 10^{7}$. Higher energy X-ray and gamma-ray measurements can further probe the high $\gamma$ $e^{\pm}$ population (which cools rapidly). If magnetic reconnection is the primary acceleration mechanism (as suggested by the very hard observed spectra), the maximum electron energy should be $\gamma_{\rm max}\sim\sigma\gamma_{\rm min}$, where $\sigma=B^{2}/(4\pi\gamma_{\rm min}n_{0}m_{e}c^{2})$ is the pulsar wind magnetization and $n_{0}$ is the e± number density (Sironi & Spitkovsky, 2011, 2014). As expected, the J2215 pulsar wind should be very strongly magnetized with $10\lesssim\sigma\lesssim 6\times 10^{7}$. Finally, the companion hot spot indicates extra localized heating. Sanchez & Romani (2017) attribute companion hot spots to IBS-energized particles precipitating to the companion magnetic poles. In our fits, the small $\beta$ allows the relatively flat IBS shock to capture $\sim 40-50\%$ of the $\dot{E}_{pw}$ flux, while the thermal emission of the hot spot represents $1\times 10^{32}$ ${\rm erg\,s^{-1}}$. Only a small fraction $\sim 2\times 10^{-3}$ of the maximum available IBS-processed pulsar wind power needs to reach the surface to heat the companion magnetic pole. Simultaneous optical/UV observations would be useful to re-check the U band normalization and further constrain the hot spot contribution and atmosphere metallicity.
J2215’s high inferred mass-loss rate makes this system a plausible isolated MSP progenitor, while its unusually flat IBS may reprocess a large fraction of the spin-down power. The hard X-ray spectrum and sharp peaks provide an important probe of relativistic shock dynamics and particle acceleration. Higher energy observations (e.g. with NuSTAR) should further constrain the IBS $e^{\pm}$ population. Since the prominent IBS peaks indicate beamed emission and thus moderate bulk Lorentz factors in the IBS flow, the various orbital phases probe different portions of the shock with varying obliquity. Thus, IBS-dominated spider pulsars, like J2215, show emission from a range of relativistic shock geometries. Future phase-resolved spectrum and polarization measurements of these natural laboratories can be compared with numerical simulations to advance our understanding of oblique, strongly magnetized relativistic shocks.